OpenStudy (anonymous):

what is a solution of the equation y=2x+5 ? ( 1point ) A. (5,20) B.(6,12) C.(3,11) D.(4,14)

OpenStudy (anonymous):

@wio @Destinymasha

jimthompson5910 (jim_thompson5910):

Let's check point A (x,y) = (5,20) so x = 5 and y = 20 y=2x+5 20 = 2*(5) + 5 ... replace x with 5, y with 20 20 = 10 + 5 20 = 15 Since 20 = 15 is a false equation, this means (2,5) is NOT a solution to y=2x+5

jimthompson5910 (jim_thompson5910):

Do the same with the other points

OpenStudy (anonymous):

so whats the answer ? im going for b or c im not sure

jimthompson5910 (jim_thompson5910):

check and show me what you get